OpenStudy (anonymous):

Find three possible rearrangements of the equation f(x)=0 into the form x=F(x). B) \[f(x) = e^x - {5 \over x}\]

OpenStudy (anonymous):

Iterations

sam (.sam.):

@satellite73 @zarkon @amistre64 @mylinaya

OpenStudy (anonymous):

Thanks .Sam.

OpenStudy (anonymous):

as in newton ralphson?

OpenStudy (anonymous):

oh i think i see start with \[e^x-\frac{5}{x}=0\] \[e^x=\frac{5}{x}\] \[x=5e^{-x}\]

OpenStudy (anonymous):

are there really 3 different forms of this? let me know and i will learn something

OpenStudy (anonymous):

as in eg. f(x) = x^3 - 2x + 3 becomes x= cube root of {2x - 3} or X^3 = 2x-3 or (2x-3)^1/3

OpenStudy (anonymous):

sorry, the last is \[x= (2x-3)^{1/3}\]

OpenStudy (anonymous):

only one of those looks like \[x=F(x)\] namely the first one

OpenStudy (anonymous):

oh i think i get it. now i need to think

OpenStudy (anonymous):

ok lets try taking the log

OpenStudy (anonymous):

\[e^x=\frac{5}{x}\] \[x=\ln(\frac{5}{x})\]

OpenStudy (anonymous):

or if you prefer \[x=\ln(5)-\ln(x)\] how about that one?

OpenStudy (anonymous):

The books answers are: \[x=5e^{-x}\] \[x= \ln5-lnx\] \[x=\sqrt {5xe^{-x}}\]

OpenStudy (anonymous):

ok two out of three isn't bad...

OpenStudy (anonymous):

No, but how did you get them?

OpenStudy (anonymous):

i started with \[e^x=\frac{5}{x}\] and for the first one i solved for x on the right hand side \[e^x=\frac{5}{x}\] \[\frac{1}{e^x}=\frac{x}{5}\] \[\frac{5}{e^x}=x\] \[5e^{-x}=x\]

OpenStudy (anonymous):

secone one i started with \[e^x=\frac{5}{x}\] and solved for the x on the left hand side \[e^x=\frac{5}{x}\] \[x=\ln(\frac{5}{x})\] \[x=\ln(5)-\ln(x)\]

OpenStudy (anonymous):

i have to think about how to get the third one

OpenStudy (anonymous):

i have to say the third one looks artificial, by which i guess i mean too complicated. we can get there by algebra, but i am not sure why you would want to do so

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Is it possible you help me with the new one I wrote up?